# CppSATD: A Reusable Self-Admitted Technical Debt Dataset in C++

### v1

@@ Abstract
1 Background & Summary
1.1 Background
Technical Debt (TD) manifests when developers have to choose quick, convenient solutions instead of more thorough and optimal ones [8]. Usually, development teams work under pressure due to user demands, with limited resources and time-to-market, practitioners might intentionally, or unintentionally accumulate TD, which can lead to increased maintenance costs, poor system performance, and an overall degradation in code quality. Self-Admitted Technical Debt (SATD) is a sub-type of TD that represents those SATD instances purposely admitted by the developers for future improvements, and typically reported via source code comments, commit messages, pull requests, or issue records. Based on the described characteristics of the issues according to the practitioners and their locations, we can categorize SATD into various types (Alves et al. [2]). Specifically, Alves et al. [2] proposed 13 SATD types differentiated by their characteristics in software systems, including architecture, build, code, defect, design, documentation, infrastructure, people, process, requirement, service, test automation, and test debt.
da Silva Maldonado et al. [9] identified SATD from over 62,000 comments of 10 Java projects and classified them into five SATD types (i.e., design, requirement, defect, test, and documentation), which are the only SATD types that can commonly be found in code comments (Alves et al. [2]). Such dataset has been widely employed in the literature (Sheikhaei et al. [26], Santos et al. [24], Chen et al. [5], Yu et al. [33]), and is commonly referred to as Maldonado-62k dataset. However, current research efforts dominantly focus on Java, revealing a gap of cross-language datasets. Moreover, the Maldonado-62k dataset does not include surrounding code contexts, which is an important component that might improve the overall accuracy of detection models. Thus, we introduce CppSATD, which is the first C++ SATD dataset with multiclass SATD annotations and surrounding code contexts, thereby allowing researchers to broaden their scope to another popular programming language and open a new avenue of research for SATD types.
In this study, we combine design and code debt as one category due to their high similarity in both characteristics and definitions, resulting in five main SATD categories: Design/Code, Requirement, Defect, Test, and Documentation. Table 1 presents the definitions and indicators of each SATD type, according to Alves et al. [2] and de Freitas Farias et al. [10].
Table 1: Definitions of SATD types. Adapted from Alves et al. [2] and de Freitas Farias et al. [10].
@@ 1.1 Background
1.2 Related Work
Prior research has contributed various datasets for the landscape of SATD studies, each with distinct characteristics and limitations. The Maldonado-62k dataset [9] and the OBrien dataset [21] captured source code comments from Java repositories and Machine Learning repositories written in Python. While these datasets attempted to classify comments into different types of SATD, they omitted crucial surrounding code contexts that could support the interpretation of developers’ intentions expressed in the comments. In contrast, the PENTACET dataset [28] mined Java comments along with their surrounding code contexts. However, PENTACET was limited to binary classification, identifying only SATD and NON-SATD instead of specific SATD types.
Conversely, Li et al. [17] combined multiple data sources, including source code comments, commit messages, pull requests, and issue tracking systems, to broaden the scope of SATD detection to other software artifacts besides source code comments. Xiao et al. [32] focused on building systems, analyzing Maven repository comments to classify them based on the locations of SATD, the underlying reasons for SATD, and the purposes that SATD stands for. On the other hand, SonarQube, Lenarduzzi et al. [15] published a dataset concentrated on detailed information of TD (e.g., commit messages, authors, time of introduction) that was extracted from all commits of 33 Apache Software Foundation projects, serving as an important resource for time series and explanatory research.
Our study aims to fill in the gaps in the existing datasets, particularly the lack of multiclass classification and the limited attention to a programming language other than Java. By annotating the comments as SATD types or NON-SATD, we present a dataset covering more than 531,000 annotated C++ source code comments with their surrounding code contexts. This vast dataset provides the research community with a novel resource for studying SATD in C++ software systems, encouraging further outcomes in this widely used language. Table 2 and Table 3 summarize the target language, the availability of contextual information, and the core software artifact of the SATD datasets introduced in prior literature.
Table 2: SATD - Binary classification datasets
References Language/Tool Code context Artifact
Sridharan et al. [28] Java Yes Comments
Xiao et al. [32] Maven No Comments
Flisar and Podgorelec [14] Java No Comments
Table 3: SATD - Multiclass classification datasets
Ref. Language/Tool Code context Artifact
da Silva Maldonado et al. [9], Bavota and Russo [3] Java No Comments
OBrien et al. [21] Python No Comments
Li et al. [17] Apache No Commit messages, Pull requests
Li et al. [16] Jira and Google Monorail issue trackers No Issues
Sharma et al. [25], Vidoni [30] R No Comments
Liu et al. [18] Python, C++, Java No Comments
1.3 Summary
With the limitations identified in prior works and the current state of SATD detection research, future studies can use CppSATD for purposes including, but not limited to:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Understanding the evolution of SATD types in C++: Tracing the lifecycle of TD in software systems has been a crucial area that was studied by Potdar and Shihab [23], Digkas et al. [11] and Sutoyo et al. [29], as it can showcase how SATD evolves in a project. Particularly, with the CppSATD dataset, the researchers will be able to examine the lifecycle of a SATD type in C++ software systems and uncover why a TD is introduced, when SATD instances are documented, how long it takes to remove them, and which type of SATD is most time-consuming to be resolved. The findings from these time-associated analyses will allow practitioners to develop corresponding strategies for each SATD type to better address the related issues, mitigate TD earlier in a software development life cycle (SDLC), and enhance long-term software maintainability.
Summary CppSATD is the first SATD comments dataset that covers C++ projects, which is one of the most widely-used programming languages [4]. Moreover, along with multi-class SATD comments, CppSATD also captures their source code contexts, providing important contextual information for each SATD instance. CppSATD is the largest available multi-type SATD dataset with surrounding code contexts. It includes over 531,000 annotated comments with 13,069 comments identified as SATD and manually classified into different SATD types (i.e., Design/Code, Requirement, Defect, Test, and Documentation).
2 Methods
@@
In this section, we present our methodology, outlining the systematic process we followed to construct the CppSATD dataset. We design our study workflow in line with the Business Process Model and Notation (BPMN) 2.0 [20] (Figure 1), and divided it into five main steps: (1) repository selection, (2) data extraction, (3) data preparation, (4) data annotation, and (5) data validation. We cover the details of the first four steps in this section, while Section 3 describes further how we performed the last step of data validation. For a quick overview, according to Figure 1, we selected the most suitable repositories based on three criteria (popularity, activeness, and diversity of domains), which resulted in five repositories: TensorFlow, React Native, Godot, Bitcoin, and Swift (Step 1). From each repository, we extracted C++ class files (.cpp and .cc files) and header files (.h files) to collect the source code comments. Using the comment locations, we retrieved their preceding and succeeding code snippets (Step 2). In step 3, we leveraged the SATD patterns, including both Easy-To-Find (ETF) and Hard-To-Find (HTF) patterns from the work of Sridharan et al. [28] to extract the comments that were most likely to be SATD for our manual annotations. The remaining comments that did not match the SATD patterns were conjectured to be NON-SATD. To assess such a conjecture, we randomly chose a sample with its size calculated to ensure a 99% confidence level and a 1% margin of error. This step generated two data subsets: a candidate SATD set (likely SATD comments) and a candidate NON-SATD set (comments with no evident SATD presence). In step 4, the authors manually annotated both subsets: the first author worked on the candidate SATD comments, while the second author was responsible for the candidate NON-SATD comments. Finally, we validated the correctness of our manual annotations by measuring the Inter-Rater Agreement (IRA) score to ensure that the annotations were consistent and reliable without any bias (Step 5). The details of each step are described in the following sections.
@@ 2.3 Data Preparation
In step 3 (Figure 1), since manually classifying over 531k comments would be an error-prone task, we filtered out the comments that were most likely to be SATD at our best efforts with a pattern-based filtering approach, and only manually annotated this subset, making the process more efficient. To perform this filtering step, we leveraged the Easy-to-Find (ETF) SATD patterns and Hard-to-Find (HTF) SATD patterns proposed in the work of Sridharan et al. [28]. In particular, ETF SATD refers to instances of SATD that can be easily spotted using simple keywords or patterns, such as ‘todo’, ‘fixme’, ‘revisit’, etc. On the other hand, HTF SATD refers to more complicated features such as ‘should be unnecessary’, ‘need[s]* to be improved’, ‘cheat here’, etc. Overall, we used 21 ETF SATD text patterns and 656 HTF text patterns to detect the most likely SATD comments. This best-effort filtering step resulted in 13,739 ETF SATD comments and 5,234 HTF SATD comments, adding up to a total of 18,973 candidate SATD comments for our manual annotation (Table 5).
@@ 2.4 Data Annotation
In particular, the first author manually classified 18,973 comments with their code contexts, resulting in 13,044 SATD comments and 5,929 NON-SATD comments (Table 7). Within the five SATD types, design/code and requirement debt are the two most prevalent categories, accounting for approximately 64% of the total comments, whereas the proportion of documentation debt is remarkably lower than other SATD types, with only 46 instances (0.24%). During the annotation process, we also encountered certain special cases where a comment could be classified into two or more types of SATD. With those scenarios, we decided to choose the category that reflected the core problem to be addressed, which was an approach followed by Maldonado and Shihab [19]. For example, consider the following SATD comment with its code context, although the SATD comment points out a test inconsistency between the two environments - which can be interpreted as test debt, the underlying issue causing the inconsistency and test failure is a defect of the system, and thus, it should be classified as defect debt.
@@
Regarding the NON-SATD random sample of 16,125 comments, the second author began the manual annotation. According to Table 8, the second author was able to identify 25 SATD comments from a total of 16,125 comments. This means that the probability of SATD instances is 0.155% in the collection of 512,394 NON-SATD comments. Hence, we can conclude with a confidence level of 99% that the remaining set of 512,394 NON-SATD comments has less than 1% chance of containing some SATD comments. Accordingly, we applied the NON-SATD label to all leftover comments after ETF and HTF pattern filtering, thereby finalizing the data annotation.
@@
3 Technical Validation
3.1 Data Validation
To evaluate the soundness of our manual annotations, we calculated the Inter-Rater Agreement (IRA) using Cohen’s kappa coefficient introduced by Cohen [6]. In particular, Cohen’s kappa coefficient evaluates the agreement and provides us with a statistical mean that can assess whether the agreement is not observed by chance, making it more accurate than other simple percent agreement calculations. The coefficient ranges from -1 to +1, and the closer the value is to +1, the more solid the agreement, which is formulated as follows:
@@ 3.1 Data Validation
To compute Cohen’s kappa coefficient, we randomly selected two representative samples for the two manually annotated data subsets. We determined the sampling sizes based on a 95% confidence level and a 5% margin of error, following Equation 1 and the methodology adopted by Sridharan et al. [28], Falessi et al. [13]. This yielded sample sizes of 374 for the 13,044 SATD comments identified by the manual annotation of the first author, and 376 for the set of 16,125 candidate NON-SATD comments annotated by the second author. To promote consistency, we scaled both to 385. We stratified the SATD sample by SATD type (including both ETF and HTF SATD), and the NON-SATD sample by SATD/NON-SATD label.
Once the two representative samples were all set, the two authors independently reviewed them for the IRA calculation. Consequently, the SATD annotation sample resulted in a Cohen’s kappa coefficient of 0.86, while the NON-SATD annotation sample gave a coefficient of 1.0. According to Sim and Wright [27], kappa scores greater than 0.80 are considered excellent (Table 9). These high IRA scores reveal excellent consistency, validating the reliability of our manual classifications [12].
@@
3.2 Threats to Validity
This section presents the limitations and threats to the validity of our dataset according to Wohlin et al. [31].
4 Data Records
@@
Code Availability
The code and the patterns used for the data extraction are available and accessible in our replication package at Zenodo [22].
Competing Interests
The authors declare that they have no competing interests, financial or otherwise, that could have appeared to influence the work reported in this paper.
Acknowledgements
Nothing to declare.

### v2

# Descriptor: C++ Self-Admitted Technical Debt Dataset (CppSATD)

@@ Abstract
[…]
IEEE SOCIETY/COUNCIL Computer Society (CS)
[…]
DATA TYPE/LOCATION C++ source code and code comments; Oulu, Finland
{IEEEkeywords}
Technical Debt, SATD, C++, Dataset, Mining Software Repositories
BACKGROUND
Technical debt (TD) manifests when developers have to choose quick, convenient solutions instead of more thorough and optimal ones [1]. Development teams usually work under pressure due to user demands, with limited resources and time-to-market [2]. Hence, practitioners might intentionally or unintentionally accumulate TD, which can lead to increased maintenance costs, poor system performance, and an overall degradation in code quality [3]. Self-Admitted Technical Debt (SATD) is a sub-type of TD that represents SATD instances purposely admitted by the developers for future improvements, and typically reported via source code comments, commit messages, pull requests, or issue records. Based on the described characteristics of the issues according to the practitioners and their locations, we can categorize SATD into various types [4]. Specifically, [4] proposed 13 SATD types differentiated by their characteristics in software systems, including architecture, build, code, defect, design, documentation, infrastructure, people, process, requirement, service, test automation, and test debt.
[5] identified SATD from over 62,000 comments of 10 Java projects and classified them into five SATD types (i.e., design, requirement, defect, test, and documentation), which are the only SATD types that can commonly be found in code comments [4]. Such dataset has been widely employed in the literature [6, 7, 8, 9], and is commonly referred to as Maldonado-62k dataset. However, current research efforts dominantly focus on Java, revealing a gap of cross-language datasets. Moreover, the Maldonado-62k dataset does not include surrounding code contexts, which is an important component that might improve the overall accuracy of detection models. Thus, we introduce CppSATD, which is the first C++ SATD dataset with multiclass SATD annotations and surrounding code contexts, thereby allowing researchers to broaden their scope to another popular programming language and open a new avenue of research for SATD types.
In this study, we combine design and code debt as one category due to their high similarity in both characteristics and definitions, resulting in five main SATD categories: Design/Code, Requirement, Defect, Test, and Documentation. Table 1 presents the definitions and indicators of each SATD type, according to [4] and [10].
Table 1: Definitions of SATD types. Adapted from [4] and [10].
@@
COLLECTION METHODS AND DESIGN
@@
In this section, we present our methodology, outlining the systematic process we followed to construct the CppSATD dataset. We design our study workflow in line with the Business Process Model and Notation (BPMN) 2.0 [11] (Figure 1), and divided it into five main steps: (1) repository selection, (2) data extraction, (3) data preparation, (4) data annotation, and (5) data validation. We cover the details of the first four steps in this section, while the next section describes further how we performed the last step of data validation. For a quick overview, according to Figure 1, we selected the most suitable repositories based on three criteria (popularity, activeness, and diversity of domains), which resulted in five repositories: TensorFlow, React Native, Godot, Bitcoin, and Swift (Step 1). From each repository, we extracted C++ class files (.cpp and .cc files) and header files (.h files) to collect the source code comments. Using the comment locations, we retrieved their preceding and succeeding code snippets (Step 2). In step 3, we leveraged the SATD patterns, including both Easy-To-Find (ETF) and Hard-To-Find (HTF) patterns from the work of [12] to extract the comments that were most likely to be SATD for our manual annotations. The remaining comments that did not match the SATD patterns were conjectured to be NON-SATD. To assess such a conjecture, we randomly chose a sample with its size calculated to ensure a 99% confidence level and a 1% margin of error. This step generated two data subsets: a candidate SATD set (likely SATD comments) and a candidate NON-SATD set (comments with no evident SATD presence). In step 4, the authors manually annotated both subsets: the first author worked on the candidate SATD comments, while the second author was responsible for the candidate NON-SATD comments. Finally, we validated the correctness of our manual annotations by measuring the Inter-Rater Agreement (IRA) score to ensure that the annotations were consistent and reliable without any bias (Step 5). The details of each step are described in the following sections.
@@ 0.3 Data Preparation
In step 3 (Figure 1), since manually classifying over 531k comments would be an error-prone task, we filtered out the comments that were most likely to be SATD at our best efforts with a pattern-based filtering approach, and only manually annotated this subset, making the process more efficient. To perform this filtering step, we leveraged the Easy-to-Find (ETF) SATD patterns and Hard-to-Find (HTF) SATD patterns proposed in the work of [12]. In particular, ETF SATD refers to instances of SATD that can be easily spotted using simple keywords or patterns, such as ‘todo’, ‘fixme’, ‘revisit’, etc. On the other hand, HTF SATD refers to more complicated features such as ‘should be unnecessary’, ‘need[s]* to be improved’, ‘cheat here’, etc. Overall, we used 21 ETF SATD text patterns and 656 HTF text patterns to detect the most likely SATD comments. This best-effort filtering step resulted in 13,739 ETF SATD comments and 5,234 HTF SATD comments, adding up to a total of 18,973 candidate SATD comments for our manual annotation (Table 3).
@@ 0.4 Data Annotation
In particular, the first author manually classified 18,973 comments with their code contexts, resulting in 13,044 SATD comments and 5,929 NON-SATD comments (Table 5). Within the five SATD types, design/code and requirement debt are the two most prevalent categories, accounting for approximately 64% of the total comments, whereas the proportion of documentation debt is remarkably lower than other SATD types, with only 46 instances (0.24%). During the annotation process, we also encountered certain special cases where a comment could be classified into two or more types of SATD. With those scenarios, we decided to choose the category that reflected the core problem to be addressed, which was an approach followed by [15]. For example, consider the following SATD comment with its code context, although the SATD comment points out a test inconsistency between the two environments - which can be interpreted as test debt, the underlying issue causing the inconsistency and test failure is a defect of the system, and thus, it should be classified as defect debt.
@@
Regarding the NON-SATD random sample of 16,125 comments, the second author was able to identify 25 SATD comments from a total of 16,125 comments (Table 6). This means that the probability of SATD instances is 0.155% in the collection of 512,394 NON-SATD comments. Hence, we can conclude with a confidence level of 99% that the remaining set of 512,394 NON-SATD comments has less than 1% chance of containing some SATD comments. Accordingly, we applied the NON-SATD label to all leftover comments after ETF and HTF pattern filtering, thereby finalizing the data annotation.
@@
VALIDATION AND QUALITY
To evaluate the soundness of our manual annotations, we calculated the Inter-Rater Agreement (IRA) using Cohen’s kappa coefficient introduced by [16]. In particular, Cohen’s kappa coefficient evaluates the agreement and provides us with a statistical mean that can assess whether the agreement is not observed by chance, making it more accurate than other simple percent agreement calculations. The coefficient ranges from -1 to +1, and the closer the value is to +1, the more solid the agreement, which is formulated as follows:
@@
To compute Cohen’s kappa coefficient, we randomly selected two representative samples for the two manually annotated data subsets. We determined the sampling sizes based on a 95% confidence level and a 5% margin of error, following Equation 1 and the methodology adopted by [12, 17, 18]. This yielded sample sizes of 374 for the 13,044 SATD comments identified by the manual annotation of the first author, and 376 for the set of 16,125 candidate NON-SATD comments annotated by the second author. To promote consistency, we scaled both to 385. We stratified the SATD sample by SATD type (including both ETF and HTF SATD), and the NON-SATD sample by SATD/NON-SATD label.
Once the two representative samples were all set, the two authors independently reviewed them for the IRA calculation. Consequently, the SATD annotation sample resulted in a Cohen’s kappa coefficient of 0.86, while the NON-SATD annotation sample gave a coefficient of 1.0. According to [19], kappa scores greater than 0.80 are considered excellent (Table 7). These high IRA scores reveal excellent consistency, validating the reliability of our manual classifications [20, 21].
@@
RECORDS AND STORAGE
@@
INSIGHTS AND NOTES
Future studies can use CppSATD for purposes including, but not limited to:
  • Understanding the evolution of SATD types in C++: Tracing the lifecycle of TD in software systems has been a crucial area that was studied by [23, 24] and [25], as it can showcase how SATD evolves in a project. Particularly, with the CppSATD dataset, the researchers will be able to examine the lifecycle of a SATD type in C++ software systems and uncover why a TD is introduced, when SATD instances are documented, how long it takes to remove them, and which type of SATD is most time-consuming to be resolved. The findings from these time-associated analyses will allow practitioners to develop corresponding strategies for each SATD type to better address the related issues, mitigate TD earlier in a software development life cycle (SDLC), and enhance long-term software maintainability.
Considering the potential of our dataset, we believe CppSATD can serve as a foundation for multiple promising extensions, including but not limited to the following areas:
  • Our multi-class SATD dataset will open future work in applying supervised learning models for SATD detection in C++. We can establish novel baseline benchmarks on C++ by using LLMs such as StarCode or QwenCoder fine-tuned on this dataset [26]. Moreover, we can extend the dataset by including additional metadata, such as cyclomatic complexity, lines of code (LOC), and comment length, supporting further research that requires rich contextual information.
  • By training models on the CppSATD dataset, future work can include integrating SATD detection directly into developer environments, such as creating real-time classification plugins for IDEs (e.g., VSCode, CLion). In addition, we can expose models via APIs, possibly through Hugging Face, to allow researchers and developers to easily access pre-trained SATD classifiers for integration into tools, automated analysis pipelines, or custom applications on private codebases.
  • Economic impact is also a meaningful aspect that should be explored. Utilizing the dataset along with temporal metadata, we can analyze the cost trade-offs between different stages of SATD life cycle and estimate the return on investment (ROI) of SATD handling strategies, creating a solid bridge between academic research and industry practices.
  • Especially, as current research landscape revolves around Java programming language, future work can include cross-language research, such as comparative analysis between C++, Python, and Java. These comparative studies can provide insights into the differences of SATD characteristics between programming languages, thereby explaining how language-specific factors can influence the occurrences of SATD in software systems.
We present the limitations and threats to the validity of our dataset as follows, according to [27].
SOURCE CODE AND SCRIPTS
The Python source code and text patterns used for our data extraction are included in the ZIP folder cppsatd.data.extraction.zip of our replication package at Zenodo [22]. We present the versions and specific details of the five selected C++ repositories in Table 2. Moreover, to convert source code to XML representations, we used srcML v1.0.0 tool [14], whose latest release was in January 2020.
ACKNOWLEDGEMENTS AND INTERESTS
This work was done independently of any grants or awards. The article authors have declared no conflicts of interest.
Phuoc Pham: Methodology, Writing - Original Draft.
Murali Sridharan: Methodology, Writing - Original Draft.
Matteo Esposito: Methodology, Writing - Review & Editing, Supervision.
Valentina Lenarduzzi: Conceptualization, Validation, Supervision Resources, Writing - Review, Project administration.